Callie is the 11-year-old daughter and dependent of Mr. and Mrs. Santo. This year, Callie filed a Form 1040 on which the only item of gross income was $10,557 interest from an investment bond portfolio that Callie inherited from a great aunt. Compute Callie’s income tax. Assume the taxable year is 2019.
Answer #1

Net unearned income = unearned income – Base amount = 10557-2200 = $8357

Taxable income = net unearned income – standard deduction = 8357-1100 = $7257

Callie’s income tax = 260+(24%*(7257-2600)) = $1378 ($1377.68)

note: if answer is required to 2 decimal places, then it will be $1377.68. kindly take note of it.